# Simplified and Generalized Masked Diffusion for Discrete Data

### v1

@@ 5 Generalization to State-dependent Masking Schedules
Given the forward conditionals and marginals, we can now compute the time reversal conditioned on $x_{0}$. The full form of $q(x_{s}|x_{t},x_{0})$ is derived in Section G.1. For $x_{t}=m$, we have
$q(x_{s}|x_{t}=m,x_{0})=q(x_{s}|x_{t}=m,{\color[rgb]{0,0,1}x_{0}},{\color[rgb]{1,0,0}x_{0}x_{0}^{\top}})=\Big{(}\frac{{\bf 1}-\alpha_{s}}{{\bf 1}-\alpha_{t}}\Big{)}^{\top}{\color[rgb]{0,0,1}x_{0}}e_{m}^{\top}x_{s}+\Big{(}\frac{\alpha_{s}-\alpha_{t}}{{\bf 1}-\alpha_{t}}\Big{)}^{\top}{\color[rgb]{1,0,0}x_{0}x_{0}^{\top}}x_{s}.$ (13)
This suggests that the backward model given $x_{t}=m$ can be chosen as $p_{\theta}(x_{s}|x_{t}=m)\triangleq q(x_{s}|x_{t}=m,{\color[rgb]{0,0,1}\mu_{\theta}(x_{t},t)},{\color[rgb]{1,0,0}\mathrm{diag}(\mu_{\theta}(x_{t},t))})$ where $\mu_{\theta}(x_{t},t)$ is a probability vector defined as in Section 3 that approximates $\mathbb{E}[x_{0}|x_{t}]$ while $\mathrm{diag}(\mu_{\theta}(x_{t},t))$ approximates $\mathbb{E}[x_{0}x_{0}^{\top}|x_{t}]=\mathrm{diag}(\mathbb{E}[x_{0}|x_{t}])$. We show in Section G.1 that the negative continuous-time ELBO for the state-dependent rate case is
@@
where $\alpha_{t}^{\prime}$ is the elementwise derivative of $\alpha_{t}$ w.r.t. $t$. This generalizes the loss from (4), which can be recovered by choosing $\alpha_{t}$ as a scalar schedule times an all-one vector. For $N$ tokens the backward model and the loss further generalize similarly to Section 3; see Section G.2.
Figure 2: Iterative unmasking process for a unconditionally generated sample by MD4. This visualization only includes a subsequence from a 1024-token sequence generated by MD4. The "?" symbol represents masked tokens. Green masked tokens are revealed in steps 500-700, followed by yellow masked tokens in steps 700-850, and finally, red masked tokens in steps 850-1000. Additional unconditional generation from our MD4 can be found in Appendix I.
@@ 6.1 Text
Text is natural discrete data with rich structures. We experiment with two datasets: text8 [41], a character-level text modeling benchmark extracted from English Wikipedia, and OpenWebText [42], an open clone of the unreleased WebText dataset used to train GPT-2 [43]. We choose the linear noise schedule described in Section 2 for all text experiments since we find it works best. We refer to our simple masked diffusion model as MD4 (Masked Discrete Diffusion for Discrete Data) and our generalized state-dependent model as GenMD4.
Figure 3: Perplexity on OpenWebText (OWT) validation set during training.
OpenWebText.
We follow Radford et al. [43], Lou et al. [32] to test the language modeling capabilities of our model. Specifically, we train a GPT2-small and GPT2-medium size of MD4 on OpenWebText (98% training, 2% validation) and evaluate zero-shot perplexity unconditionally on the test splits of five benchmark datasets used in Radford et al. [43]: LAMBADA [44], WikiText2, Penn Treebank [45], WikiText103 [46], and One Billion Words [47]. We keep our evaluation setup the same as Lou et al. [32]. To ensure the comparison is fair, we reproduced their result in our implementation.
The results are shown in LABEL:tab:owt-zeroshot-ppl. We can see that our small model outperforms previous best discrete diffusion models on all five tasks. We also confirmed that our re-implementation of SEDD Absorb leads to similar and slightly better results than those reported in their paper. We believe this is because they were using mixed-precision training (thus larger variances) while we are using full precision (float32) training. We are also better than GPT-2 on all tasks except LAMBADA where we are the second best among all methods. When we scale up the model size to medium, we observe that MD4 similarly beats SEDD Absorb and GPT-2 on all tasks except LAMBADA.
@@
To validate that the strong zero-shot performance is a result of a better-trained model, we plot perplexity on the OpenWebText validation set in Figure 3, it demonstrates that our model converges faster and has better final likelihoods than prior methods. We also observed that SEDD [32] has training instabilities. We believe this is due to the score parameterization breaking consistency between forward and reverse processes, as explained in Section 4. Although GenMD4 achieves lower perplexity than MD4, we observed that the learned $w$s can easily overfit to dataset statistics, making it less effective on zero-shot transfer tasks. Beyond likelihood evaluation, we also examine generation quality of our models. Figure 2 showcases a randomly selected, notably coherent sample produced by MD4-medium, alongside its iterative denoising process. Quantitatively, we employ GPT-2 Large model to evaluate perplexity of the generated samples as shown in Figure 6. We can see MD4 is able to outperform its auto-regressive counterpart GPT-2 by a large margin. Sample quality can be improved by two orthogonal strategies: scaling up model size (more compute during training) or increasing decoding steps (more compute during inference). Compared with SEDD, MD4 reaches a new state of the art. Notably, after 64 decoding steps, our MD4 of small size matches the performance of SEDD-medium.
@@
Following the convention in Austin et al. [14], Lou et al. [32], we experimented with training masked diffusion models on text chunks of length 256. We used the same dataset splits and transformers of the same size to parameterize the backward prediction model. Details can be found in Section H.1. Results are summarized in Table 3. We used the standard bits-per-character metric (a normalized negative log likelihood) for this dataset. We can see that our models outperform previous diffusion models, whether discrete or continuous in nature. We also outperform the best any-order autoregressive models, which do not assume a fixed generation order and has a strong connection to discrete diffusion [48]. Our model is only beaten by an autoregressive transformer and the Discrete Flow (which also used an autoregressive backbone). We believe this is because autoregressive models only require learning a fixed order of generation thus better utilize model capacity.
@@
SEDD Absorb [32] $\leq$ 1.41333Our personal communication with the authors of Lou et al. [32] concluded that the paper’s reported result (1.32) is accidentally computed on the training set. Here we report our reproduced result after correcting the mistake.
@@ 6.2 Pixel-level image modeling
Figure 4: Non cherry-picked unconditional samples from MD4 trained on ImageNet 64x64, treating pixels as discrete tokens. More samples can be found in Figure 5 in Appendix. The model is optimized for likelihood instead of visual quality—see e.g., Kingma et al. [33] for samples from a continuous diffusion model optimized similarly for likelihood.
@@
Results are summarized in Table 3. We establish a new state-of-the-art for discrete diffusion models, outperforming previous discrete-time [14] and continuous-time [29] models by a significant margin. Notably, despite lacking knowledge of the ordinal structure of pixel values, MD4 still outperforms models trained with this inductive bias, including D3PM Gauss and $\tau$LDR where the noising distribution is a discrete Gaussian distribution that gives larger probabilities to near pixel values. Our CIFAR-10 model without data augmentation also achieves better likelihood than the previous best autoregressive models on this task, while on ImageNet $64\times 64$ our result is competitive with Transformer AR models of similar sizes, only beaten by Perceiver AR which is approximately 4×4$\times$4 × larger in model size. We provide a random sample from our ImageNet 64$\times$64 model in Figure 4. More results can be found in in Section H.3.
@@ 7 Conclusion
In this work, we revisit masked diffusion models, focusing on a flexible continuous-time formulation. Existing works in this area are not easily accessible to non-specialists and present ELBOs that are difficult to optimize, often resulting in performance that is not competitive with continuous diffusions and autoregressive models. The framework we propose provides a very simple expression of the ELBO as a weighted integral of cross-entropy losses. Additionally, we propose a generalized masked diffusion formulation (GenMD4), where the masking schedule depends on the current state of the process, and derive its corresponding ELBO. Experimentally, on text and image data, the resulting masked diffusions outperform existing discrete and continuous diffusion models and fare very well compared to autoregressive models for most tasks. GenMD4 provides further improvements in terms of likelihoods over the state-independent case.
@@ Appendix C Time reversal of the forward process given $x_{0}$
We can rewrite the above using reverse transition matrix $\bar{R}^{x_{0}}(t,s)\in\mathbb{R}^{(m+1)\times(m+1)}$ as
@@
We are also interested in what would happen in the infinitesimal time limit, i.e., when $s=t-\Delta t$ and $\Delta t\to 0$. Note that $\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}=-\beta(t)\alpha_{t}$,
$\alpha_{t-\Delta t}-\alpha_{t}=\alpha_{t}^{\prime}\Delta t+o(\Delta t).$
@@
$\bar{R}^{x_{0}}(t,t-\Delta t)=I+\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}e_{m}(x_{0}-e_{m})^{\top}\Delta t+o(\Delta t).$
@@
$R^{x_{0}}(t)$ $=\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}e_{m}(x_{0}-e_{m})^{\top}.$ (20)
@@ Appendix D Details of the ELBO
Using (C) and (3), we compute the KL divergences between forward and backward transitions
@@
Note that $00\log 0=$ 0. In this case, the reconstruction term becomes
@@ Appendix E Avoiding undefined KL divergence
Appendix F Unifying Existing Masked Diffusion Models
F.1 The CTMC point of view

### v2

@@ 6 Generalization to State-dependent Masking Schedules
Given the forward conditionals and marginals, we can now compute the time reversal conditioned on $x_{0}$. The full form of $q(x_{s}|x_{t},x_{0})$ is derived in Sec. I.1. For $x_{t}=m$, we have
$q(x_{s}|x_{t}=m,x_{0})=q(x_{s}|x_{t}=m,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{0}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}x_{0}^{\top}})=\Big{(}\frac{{\bf 1}-\alpha_{s}}{{\bf 1}-\alpha_{t}}\Big{)}^{\top}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}x_{0}}e_{m}^{\top}x_{s}+\Big{(}\frac{\alpha_{s}-\alpha_{t}}{{\bf 1}-\alpha_{t}}\Big{)}^{\top}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}x_{0}^{\top}}x_{s}.$ (11)
This suggests that the reverse model given $x_{t}=m$ can be chosen as $p_{\theta}(x_{s}|x_{t}=m)\triangleq q(x_{s}|x_{t}=m,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mu_{\theta}(x_{t},t)},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{diag}(\mu_{\theta}(x_{t},t))})$ where $\mu_{\theta}(x_{t},t)$ is a probability vector defined as in Sec. 3 that approximates $\mathbb{E}[x_{0}|x_{t}]$ while $\mathrm{diag}(\mu_{\theta}(x_{t},t))$ approximates $\mathbb{E}[x_{0}x_{0}^{\top}|x_{t}]=\mathrm{diag}(\mathbb{E}[x_{0}|x_{t}])$. We show in Sec. I.1 that the negative continuous-time ELBO for the state-dependent rate case is
@@
where $\alpha_{t}^{\prime}$ is the elementwise derivative of $\alpha_{t}$ w.r.t. $t$. This generalizes the loss from (4), which can be recovered by choosing $\alpha_{t}$ as a scalar schedule times an all-one vector. For $N$ tokens the reverse model and the loss further generalize similarly to Sec. 3; see Sec. I.2.
Figure 3: Iterative unmasking process for an unconditionally generated sample by MD4. This visualization only includes a subsequence from a generated sequence of 1024 tokens. "?" represents masks. Masked tokens are revealed sequentially: green (steps 500-700), yellow (700-850), and red (850-1000). Additional unconditional generation from MD4 can be found in Sec. K.4.
@@ 7.1 Text
Text is natural discrete data with rich structures. We experiment with two datasets: text8 [54], a character-level text modeling benchmark extracted from English Wikipedia, and OpenWebText [55], an open clone of the unreleased WebText dataset used to train GPT-2 [56]. We choose the linear masking schedule described in Sec. 2 for all text experiments since we find it works best. We refer to our simple masked diffusion model as MD4 (Masked Discrete Diffusion for Discrete Data) and our generalized state-dependent model as GenMD4.
@@
Figure 4: Perplexity on OpenWebText (OWT) validation set during training. The final numbers are reported in Tab. 5 in Appendix.
OpenWebText.
We train a GPT2-small and GPT2-medium size of MD4 on OpenWebText (98% training, 2% validation) and evaluate zero-shot perplexity on the test splits of five benchmark datasets used in Radford et al. [56]: LAMBADA [57], WikiText2, Penn Treebank [58], WikiText103 [59], and One Billion Words [60]. We keep our evaluation setup the same as SEDD [32]. To ensure the comparison is fair, we reproduced their result in our implementation.
The results are shown in LABEL:tab:owt-zeroshot-ppl. Our small model outperforms previous best discrete diffusion models on all five tasks. We also confirmed that our re-implementation of SEDD leads to similar and slightly better results than those reported in their paper. We believe this is because they were using mixed-precision training (thus larger variances) while we are using full precision (float32) training. We are also better than GPT-2 on all tasks except LAMBADA where we are the second best among all methods. When we scale up the model size to medium, we observe that MD4 similarly beats SEDD and GPT-2 on all tasks except LAMBADA.
To validate that the strong zero-shot performance is a result of a better-trained model, we plot perplexity on the OpenWebText validation set in Fig. 4. It demonstrates that our model converges faster and has better final likelihoods than prior methods. We also observed that SEDD [32] has training instabilities. We believe this is due to the score parameterization breaking consistency between forward and reverse processes, as explained in Sec. 5. Although GenMD4 achieves lower perplexity than MD4, we observed that the learned $w$s can easily overfit to dataset statistics, making it less effective on zero-shot transfer tasks. Beyond likelihood evaluation, we also examine generation quality of our models. Fig. 3 showcases a randomly selected, notably coherent sample produced by MD4-medium, alongside its iterative denoising process. Fig. 9 demonstrates MD4’s ability to perform text infilling and highlights a substantial improvement in sample quality when transitioning from the linear to the cosine schedule.
We follow prior work to employ the GPT-2 Large model to evaluate perplexity of the generated samples. As noted in Sec. 4, the generative perplexity metric can be misleadingly decreased by reducing sample diversity; we include these results in Appendix (Fig. 7) mainly for completeness. Under this metric, MD4 significantly outperforms the GPT-2 checkpoint and SEDD. Notably, after 64 decoding steps, our small-sized MD4 matches the performance of SEDD-medium.
@@
Following the convention in prior work [14, 32], we experimented with training masked diffusion models on text chunks of length 256. We used the same dataset splits and transformers of the same size to parameterize the prediction model. Details can be found in Sec. J.1. Results are summarized in Tab. 3. We used the standard bits-per-character metric (a normalized negative log likelihood) for this dataset. We can see that our models outperform previous diffusion models, whether discrete or continuous in nature. We also outperform the best any-order autoregressive models, which do not assume a fixed generation order and has a strong connection to discrete diffusion [47]. Our model is only beaten by an autoregressive transformer and the Discrete Flow (which also used an autoregressive backbone). We believe this is because autoregressive models only require learning a fixed order of generation thus better utilize model capacity.
@@
SEDD Absorb [32] $\leq$ 1.39
@@ 7.2 Pixel-level image modeling
Figure 5: Non cherry-picked unconditional samples from MD4 trained on ImageNet 64x64, treating pixels as discrete tokens. More samples can be found in Fig. 8 in Appendix. The model is optimized for likelihood instead of visual quality—see e.g., Kingma et al. [33] for samples from a continuous diffusion model optimized similarly for likelihood.
@@
Results are summarized in Tab. 3. We establish a new state-of-the-art for discrete diffusion models, outperforming previous discrete-time [14] and continuous-time [29] models by a significant margin. Notably, despite lacking knowledge of the ordinal structure of pixel values, MD4 still outperforms models trained with this inductive bias, including D3PM Gauss and $\tau$LDR [29] where the noising distribution is a discrete Gaussian distribution that gives larger probabilities to near pixel values. To isolate the differences caused by training objectives, we also implemented the objective from Campbell et al. [29] in our code. The result, labeled as $\tau$LDR Mask, demonstrates that the high variance of this objective negatively impacts model learning, even when using the same architecture as ours. Our CIFAR-10 model without data augmentation also achieves better likelihood than the previous best autoregressive models on this task, while on ImageNet $64\times 64$ our result is competitive with Transformer AR models which is approximately 4×4$\times$4 × larger in model size and the continuous diffusion model VDM [33].
We provide a random sample from our ImageNet 64$\times$64 model in Fig. 5. More results can be found in in App. K. In Fig. 2, we plot the FID values of samples generated under different choices of schedules and discretization grids. We can see that the model with the linear schedule and a cosine grid achieves an FID close to the model with cosine schedule, both significantly outperform the linear schedule with a uniform grid. We further trained a class-conditional model on ImageNet 64$\times$64 that boosts the FID to around 7. Although these are not state-of-the-art FIDs on ImageNet 64$\times$64, we emphasize our models are optimized for likelihood instead of sample quality.
@@ 8 Conclusion
In this work, we revisit masked diffusion models, focusing on a flexible continuous-time formulation. Existing works in this area are not easily accessible to non-specialists and present ELBOs that are difficult to optimize, often resulting in performance that is not competitive with continuous diffusions and autoregressive models. The framework we propose provides a very simple expression of the ELBO as a weighted integral of cross-entropy losses. Additionally, we propose a generalized masked diffusion formulation (GenMD4), where the masking schedule depends on the current state of the process, and derive its corresponding ELBO. On text data, our MD4 models outperform existing discrete and continuous diffusion models. For pixel-level image modeling, we significantly improve discrete diffusion results, outperforming similar-sized AR models and achieving comparable likelihoods to continuous diffusion models such as VDM. GenMD4 provides further improvements in terms of likelihoods over the state-independent case.
@@ Appendix C Time reversal of the forward process given $x_{0}$
Visually, eqn (C) is a $\mathbb{R}^{(m+1)\times(m+1)}$ matrix (Fig. 6, left) whose first index is $x_{t}$ and the second is $x_{s}$. The matrix is almost an identity matrix except the last row corresponding to $x_{t}$ is the mask token. The last row means with probability of $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$ the mask token gets unmasked to become $x_{0}$, and with probability of $\frac{1-\alpha_{s}}{1-\alpha_{t}}$ it remains masked.
Alternatively, we can rewrite the above using reverse transition matrix $\bar{R}^{x_{0}}(t,s)\in\mathbb{R}^{(m+1)\times(m+1)}$ as
@@
Figure 6: The reverse transition probability and our generative model. Left: $q(x_{s}=\cdot|x_{t}=\cdot,x_{0})$ in matrix form where first index is $x_{t}$ and second index is $x_{s}$. Right: $p_{\theta}(x_{s}=\cdot|x_{t}=\cdot)\triangleq q(x_{s}=\cdot|x_{t}=\cdot,\mu_{\theta}(x_{t},t))$ also in matrix form.
We are also interested in what would happen in the infinitesimal time limit, i.e., when $s=t-\Delta t$ and $\Delta t\to 0$. Note that
$\alpha_{t-\Delta t}-\alpha_{t}=-\alpha_{t}^{\prime}\Delta t+o(\Delta t).$
@@
$\bar{R}^{x_{0}}(t,t-\Delta t)=I-\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}e_{m}(x_{0}-e_{m})^{\top}\Delta t+o(\Delta t).$
@@
$R^{x_{0}}(t)$ $=-\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}e_{m}(x_{0}-e_{m})^{\top}.$ (18)
@@ Appendix D Details of the ELBO
Using (C) and (3), we compute the KL divergences between forward and reverse transitions
@@
Note that $00\log 0=$ 0. Alternatively, this result can be easily obtained from the visual depictions of $q(x_{s}|x_{t},x_{0})$ and $p_{\theta}(x_{s}|x_{t})$ shown in Fig. 6. In this case, the reconstruction term becomes
@@ Appendix E Avoiding undefined KL divergence
Appendix F Details of Training and Sampling with MD4
F.1 Training
Algorithm 1 A single step of training with MD4.
Input: data minibatch $\{x_{t}^{i}\}_{i=1}^{B}$, network $\mu_{\theta}(\cdot,t)$, masking schedule $\alpha_{t}$
for $i=1,\dots,B$ do (in parallel):
     $t_{i}\leftarrow\mathrm{mod}(u+i/B,1)$, $u\sim U[0,1]$
     for $n\in[N]$, mask out each token $x_{0}^{i,(n)}$ independently with probability $1-\alpha_{t_{i}}$ to obtain $x_{t_{i}}^{i}$
     for $\text{for }n\in[N]$, if $x_{t_{i}}^{(n)}\!=\!m$, compute weighted cross entropy loss $\frac{\alpha_{t_{i}}^{\prime}}{1-\alpha_{t_{i}}}(x_{0}^{i,(n)})^{\top}\log\mu^{(n)}_{\theta}(x_{t_{i}}^{i},t_{i})$
Sum over all weighted cross entropy losses for mask positions and optimize via autodiff
F.2 Sampling
Algorithm 2 Unconditional and conditional generation (e.g., infilling) with MD4.
Input: Context sequence $x^{c}$ of length $N$, with masks indicating the target areas for generation
Init: $\{t(i)\}_{i=0}^{T}\leftarrow\text{discretize}([0,1])$, $x_{t(T)}\leftarrow x^{c}$
for $i=T,T-1,\dots,1$ do
     $t\leftarrow t(i)$, $s\leftarrow t(i-1)$
     for $\text{for }n\in[N]$, if $x_{t}^{(n)}=m$, draw $x_{s}^{(n)}\sim\mathrm{Cat}(\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\mu_{\theta}^{(n)}(x_{t},t)+\frac{1-\alpha_{s}}{1-\alpha_{t}}e_{m})$ else $x_{s}^{(n)}\leftarrow x_{t}^{(n)}$
return $x_{0}$.
Appendix G JAX Categorical Sampling and Implicit Top-$p$
We noticed that the following equivalent implementation of Alg. 2 leads to significantly worse sample quality in JAX:
Algorithm 3 Variant of Alg. 2 that yields lower sample quality when implemented in JAX.
Input: Token sequence $x^{c}$ of length $N$, with masks indicating the target areas for generation
Init: $\{t(i)\}_{i=0}^{T}\leftarrow\text{discretize}([0,1])$, $x_{t(T)}\leftarrow x^{c}$
for $i=T,T-1,\dots,1$ do
     $t\leftarrow t(i)$, $s\leftarrow t(i-1)$
     for $n\in[N]$ do (in parallel)
         draw $u\sim U[0,1]$
         if $x_{t}^{(n)}=m$ and $u<\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$ then
              draw $x_{s}^{(n)}\sim\mathrm{Cat}(\mu_{\theta}^{(n)}(x_{t},t))$
         else
              $x_{s}^{(n)}\leftarrow x_{t}^{(n)}$
return $x_{0}$.
However, mathetically it is equivalent to Alg. 2 and should produce identical results. Our investigation revealed that the issue arises because Alg. 2 scales the output probabilities of $\mu_{\theta}$ by a small factor $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$ as $s$ is close to $t$, causing some categories to have very low probabilities. JAX, however, implements categorical sampling using Gumbel argmax, which is less numerically stable than methods like binary search. As a result, categories with low probabilities are rarely sampled, even when their cumulative probability is significant. In our experiment, we found that categories with probabilities below 1e-8 are rarely sampled out of a total of 50K categories. Thus, Alg. 2 implicitly performs top-$p$ sampling (with a dynamic p) under JAX’s categorical sampling, yielding better sample quality than Alg. 3 where $\mu_{\theta}$ is not scaled by a small factor and has fewer categories truncated.
Appendix H Unifying Existing Masked Diffusion Models
H.1 The CTMC point of view